OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (If not possible, enter IMPOSSIBLE.) 3x − 9 4x − 16

OpenStudy (anonymous):

It's a fraction.

hero (hero):

\(\dfrac{3x - 9}{4x - 16} = \dfrac{3(x - 3)}{4(x - 4)}\) \(=\dfrac{3}{4} \dot\ \dfrac{x - 3}{x - 4}\) \(=\dfrac{3}{4} \dot\ \dfrac{x - 4 + 1}{x - 4}\) \(=\dfrac{3}{4} \left( \dfrac{x - 4}{x - 4} + \dfrac{1}{x - 4}\right)\) \(=\dfrac{3}{4} \left( 1+ \dfrac{1}{x - 4}\right)\) \(=\dfrac{3}{4} + \dfrac{3}{4(x - 4)}\)

OpenStudy (anonymous):

Specify the restrictions on the variable. (Select all that apply.) x ≠ 4 x ≠ 16 x ≠ 0 x ≠ 3 x ≠ 9 (Which one would be the answer?) @Hero

hero (hero):

Which one do you think would be the answer? Remember, the denominator cannot equal zero.

OpenStudy (anonymous):

I think it was 3 & 4

hero (hero):

What do you mean 3 and 4? There's only one value that x cannot equal. Which is it?

OpenStudy (anonymous):

0? I missed school when this was taught, I have no idea, sorry.

hero (hero):

\[4x - 16 \ne 0\] Solve that to find out what x cannot equal

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Is that the correct answer? @Hero

hero (hero):

\[x \ne 4\] is correct, yes

OpenStudy (anonymous):

Ok, thank you...the next question, I think I got the answer, but can you see if it's correct for me, please?

OpenStudy (anonymous):

The answer is incorrect, I don't think the answer is all the way reduced. @ganeshie8 care to help?

ganeshie8 (ganeshie8):

which question ?

hero (hero):

No, it is, I just factored it too much

hero (hero):

The correct answer is \[\frac{3(x - 3)}{4(x - 4)}\]

OpenStudy (anonymous):

oooooh ! ok Thank you lol I think I had that but didnt understand the second part

ganeshie8 (ganeshie8):

The rational expression in question not like below, right ? \(\dfrac{1}{(3x-9)(4x-16)}\)

OpenStudy (anonymous):

no, the 3x-9 is on top.

ganeshie8 (ganeshie8):

good, then \(\dfrac{3(x - 3)}{4(x - 4)}\) is the final simplified form

ganeshie8 (ganeshie8):

this expression goes crazy when the denominator stuff equals \(0\)

ganeshie8 (ganeshie8):

So, to get the restrictions : set the denominator equal to \(0\), and solve the variable

OpenStudy (anonymous):

Ok, hopefully, I'll remember lol

ganeshie8 (ganeshie8):

:) thats one important thing to keep in mind when u deal with solving equations involving rational expressions.

OpenStudy (anonymous):

Ok, thank you...Can you help me with the other problem i tagged you in?